### v1

# The Regular property of Invariant Rings over a Dedekind Domain

Abstract.
Let AA be a Dedekind domain and let $K$ be its field of fractions. Let $G\subseteq GL_{n}(A)$ be a finite group. Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing AA). Assume that $|G|$ is invertible in AA. We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Key words and phrases:
invariant rings, ring of integers of global fields, regular rings, group cohomology.
2020 Mathematics Subject Classification:
13A50, 13H05
1. Introduction
Consider the hypothesis below.
Hypothesis $1.$
Let AA be a Dedekind domain and let $K$ be its field of fractions. Consider a finite group $G\subseteq GL_{n}(A)$. Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing AA). Assume that $|G|$ is invertible in AA.
In this article we establish the following result which is a variant of Theorem 7.2.1 in [1] for polynomial ring over a Dedekind domain.
Theorem 1.2.
Assume Hypothesis $1.$ Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
Example 1.3.
The conclusion of Theorem 1.2 is true in the following cases.
  1. (1)
    AA is the ring of integers of a global function field i.e. a finite extension of $\mathbb{F}_{q}$(t) where $q$ is some power of a prime $p$, and $p\nmid|G|$.
  2. (2)
    AA is localization at a prime of the ring of integers of a number field and $|G|$ is non-zero in the residue field.
Remark 1.3.1.
We might wonder whether Theorem 1.2 holds if AA is a regular domain of higher dimension. However our technique fails when $\dim A\geq 2$, see 2.3.1.
2. Auxiliary Results
Consider the hypothesis below.
Hypothesis 2.1.
Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is a field). Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$ with $|G|=r$ invertible in $\mathcal{O}/\pi\mathcal{O}$.
We can easily generalize Proposition 3.5 in [2] (which was for $n=2$) for any $n\in\mathbb{N}$ for our case. We provide the proof below for the sake of completeness.
Proposition $2.$
Assume Hypothesis 2.1. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is injective.
Proof.
We consider elements of $GL_{n}(\mathcal{O})$ and $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ as matrices. Let $W\in G$ with $\eta(W)=I$. Then $W=I+Y$ where all entries of $Y$ are in $\pi\mathcal{O}$. Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$. Thus
$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
Set
$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
So $YU=0$. The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is rIrI which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$. By Nakayama’s Lemma it follows that this map is surjective. Since $\mathcal{O}^{n}$ is noetherian module over $\mathcal{O}$, the map is an isomorphism. Thus $U\in GL_{n}(\mathcal{O})$. As $YU=0$, we get $Y=0$. So $\eta$ is injective.∎
The following is an integral lemma for proving Theorem 1.2.
Lemma 2.3.
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$ and any psuedo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
Proof.
We prove the result by induction on $n\geq 2$. Let $n=2$ and let $\{v_{1},v_{2}\}$ be a basis of $\mathcal{O_{\pi}}^{2}$ such that $\sigma(v_{1})=v_{1}$ and $\sigma(v_{2})=\lambda v_{2}$. We can assume $v_{1},v_{2}\in\mathcal{O}^{2}\backslash\pi\mathcal{O}^{2}$. Note $\lambda\neq 1$. Let $w_{1}=v_{1}$ and let $\{w_{1},w_{2}\}$ be a basis of $\mathcal{O}^{2}$. Let $v_{2}=\alpha w_{1}+\beta w_{2}$. Note that $\beta\neq 0$. Now $\sigma(v_{2})=\lambda v_{2}=\lambda\alpha w_{1}+\lambda\beta w_{2}$. Also we have $\sigma(v_{2})=\alpha\sigma(w_{1})+\beta\sigma(w_{2})=\alpha w_{1}+\beta\sigma(w_{2})$. Thus comparing both equations we have
(2.1) $\beta\sigma(w_{2})=(\lambda-1)\alpha w_{1}+\lambda\beta w_{2}$
Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$. Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid r$. As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$. By Proposition 2.2, the map $\lambda>\rightarrow(\mathcal{O}/\pi\mathcal{O})^{\times}$ is injective. As $\lambda\neq 1$, we have that the residue of $\lambda$ in $\mathcal{O}/\pi\mathcal{O}$ is $\bar{\lambda}\neq 1$. Thus $\lambda-1\not\in\pi\mathcal{O}$. So $\lambda-1\in\mathcal{O}^{\times}$.
We claim that $\beta\in\mathcal{O}^{\times}$. Assume on the contrary $\beta\in\pi\mathcal{O}$. Since $\alpha w_{1}+\beta w_{2}=v_{2}\not\in\pi\mathcal{O}^{2}$. Thus $\alpha\not\in\pi\mathcal{O}$. Going modulo $\pi$ in the Equation 2.1, we have $(\bar{\lambda}-1)\bar{\alpha}\bar{w_{1}}=0$ which is a contradiction. Thus we can write Equation 2.1 as $\sigma(w_{2})=\lambda w_{2}+\theta w_{1}$ for some $\theta\in\mathcal{O}$. Then letting $w^{\prime}_{2}=\frac{\theta}{\lambda-1}w_{1}+w_{2}$, we have that $\{w_{1},w^{\prime}_{2}\}$ is a basis and $\sigma(w^{\prime}_{2})=\lambda w^{\prime}_{2}$.
Let $n\geq 3$ and $\{v_{1},v_{2},\dots,v_{n}\}$ be a basis of $\mathcal{O}_{\pi}$ with $v_{i}\in\mathcal{O}^{n}\backslash\pi\mathcal{O}^{n}$ and $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$ and $\lambda\neq 1$. Let $E=\mathcal{O}v_{1}$ and $w_{1}=v_{1}$. We have the following exact sequence.
$0\rightarrow E\rightarrow W=\mathcal{O}^{n}\rightarrow\bar{W}=\mathcal{O}^{n-1}\rightarrow 0.$
Now $\sigma|_{\bar{W}}$ is also a psuedo-reflection. By induction we have a basis $\{\bar{w_{2}},\bar{w_{3}},\dots,\bar{w_{n}}\}$ of $\bar{W}$ with $\sigma(\bar{w_{i}})=\bar{w_{i}}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda\bar{w_{n}}$. Now $\{w_{1},w_{2},\dots,w_{n}\}$ is a basis of $W$ and $\sigma(w_{1})=w_{1}$ and $\sigma(w_{i})=w_{i}+a_{i}w_{1}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}+a_{n}w_{1}$. We will show that $a_{2}=a_{3}=\dots=a_{n-1}=0$. Consider $2\leq i\leq n-1$. Now $\sigma(w_{i})=w_{i}+a_{i}w_{1}$. Thus $\sigma^{j}(w_{i})=w_{i}+ja_{i}w_{1}$. Now $\sigma^{m}=id$ where $m=ord(\sigma)=ord(\lambda)$. Thus $w_{i}=w_{i}+ma_{i}w_{1}$. Hence $ma_{i}w_{1}=0$. As $m$ is invertible in $\mathcal{O}$, we have $a_{i}=0$. Also by Proposition 2.2, $\lambda-1$ is invertible in $\mathcal{O}$. Thus $w_{1},w_{2},\dots,w_{n-1},w^{\prime}_{n}$ is the required basis where $w^{\prime}_{n}=\frac{a_{n}}{\lambda-1}w_{1}+w_{n}$ and $\sigma(w^{\prime}_{n})=\lambda w^{\prime}_{n}$. ∎
Corollary 2.3.1.
Assume Hypothesis 2.1. Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$. Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by psuedo-reflections.
Proof.
Follows from Proposition 2.2 and Lemma 2.3. ∎
Remark 2.3.1.
Our proof of Lemma 2.3 fails when $\mathcal{O}$ has dimension $\geq 2$. Suppose
$\mathcal{O}=k[X,Y]_{(X,Y)}$. Let $v=e_{1}/x^{2}+e_{2}/y^{2}\in K^{2}$ be such that $\sigma(v)=\lambda v$. Then note that for no choice of $t\in\mathcal{O}$ does $tv\in\mathcal{O}^{2}\setminus\mathfrak{m}\mathcal{O}^{2}$. So our induction fails at the base case $n=2$.
3. Proof of Theorem 1.2
Firstly we observe the following.
Hypothesis 3.1.
Consider the case in Hypothesis $1.$ Let $P$ be a maximal ideal of AA. Set $(\mathcal{O},(\pi))=(A_{P},PA_{P})$. Now $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}=K$. Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(\mathcal{O})$. Further since $|G|$ is invertible in AA, we have $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$. Thus for proving Theorem 1.2, it suffices to prove that $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular. Now let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$.
Now we give the proof of Theorem 1.2 by proving $S$ is regular in Hypothesis 3.1.
Proof.
Consider the short exact sequence
$0\rightarrow R\xrightarrow{\pi}R\rightarrow R/\pi R\rightarrow 0.$
We have the long exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{\pi}H^{1}(G,R)\rightarrow H^{1}(G,R/\pi R)\rightarrow\dots$
Let $K=ker(\pi)=Img(\delta)$ in the above sequence. Now we have that $K$ is $R^{G}/\pi R^{G}$-module. By Theorem 6.5.8 in [3] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$. Hence $r\cdot K=0$. As $r$ is a unit in $R^{G}/\pi R^{G}$, we have $K=0$. Hence we have the short exact sequence
$0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\rightarrow 0.$
Thus $R^{G}/\pi R^{G}\cong(R/\pi R)^{G}$. Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field. Also by Corollary 2.3.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections. Thus by Theorem 7.2.1 in [1], $(R/\pi R)^{G}$ is regular. Let $\mathfrak{m}$ be the maximal homogenous ideal of $R^{G}$. Since $\pi\not\in\mathfrak{m}^{2}$. Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎

### v2

@@
-The Regular property of Invariant Rings over a Dedekind Domain
+The Regular property of Invariant Rings over Regular Domains
 Abstract.
-Let AA be a Dedekind domain and let $K$ be its field of fractions.
+The main result of this paper is a generalization of the theorem of Chevalley-Shephard-Todd to the rings of invariants of pseudo-reflection groups over regular domains.
+More precisely, let $A$ be a regular domain and let $K$ be its field of fractions.
 Let $G\subseteq GL_{n}(A)$ be a finite group.
-Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing AA).
-Assume that $|G|$ is invertible in AA.
+Let $G$ act linearly on $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$).
+Assume that $|G|$ is invertible in $A$.
 We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
 Key words and phrases:
-invariant rings, ring of integers of global fields, regular rings, group cohomology.
+invariant rings, regular domains, regular rings, group cohomology.
 2020 Mathematics Subject Classification:
@@
 1. Introduction
-Consider the hypothesis below.
-Hypothesis $1.$
-Let AA be a Dedekind domain and let $K$ be its field of fractions.
+We have the famous Chevalley-Shephard-Todd Theorem (Theorem 7.2.1 in [1]) which states:
+Theorem 1.1.
+Let $K$ be a field. Consider a finite group $G\subseteq GL_{n}(K)$. Consider the ring
+$K[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $K$).
+Assume that $|G|$ is invertible in $K$.
+We have $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
+Mundelius had generalized Theorem 1.1 to the rings of invariants of pseudo-reflection groups over Dedekind domains (Theorem 3.8 in [2]).
+In this paper we establish the following result which is a further generalization of Theorem 1.1 to the rings of invariants of pseudo-reflection groups over regular domains.
+Theorem 1.2.
+Let $A$ be a regular domain and let $K$ be its field of fractions.
 Consider a finite group $G\subseteq GL_{n}(A)$.
-Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing AA).
-Assume that $|G|$ is invertible in AA.
-In this article we establish the following result which is a variant of Theorem 7.2.1 in [1] for polynomial ring over a Dedekind domain.
-Theorem 1.2.
-Assume Hypothesis $1.$ Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
-Example 1.3.
-The conclusion of Theorem 1.2 is true in the following cases.
-  1. (1)
-    AA is the ring of integers of a global function field i.e. a finite extension of $\mathbb{F}_{q}$(t) where $q$ is some power of a prime $p$, and $p\nmid|G|$.
-  2. (2)
-    AA is localization at a prime of the ring of integers of a number field and $|G|$ is non-zero in the residue field.
-Remark 1.3.1.
-We might wonder whether Theorem 1.2 holds if AA is a regular domain of higher dimension. However our technique fails when $\dim A\geq 2$, see 2.3.1.
-2. Auxiliary Results
+Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$).
+Assume that $|G|$ is invertible in $A$.
+If $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections, then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+In Section 2, we prove Theorem 1.2 for the case when $A$ is a DVR as Theorem 2.2.
+This also leads to an alternate proof for the case of Dedekind domains considered by Mundelius in [2].
+In Section 3, we prove Theorem 1.2 for the case when $A$ is a regular local ring as Theorem 3.2.
+Finally in Section 4, we complete the proof of Theorem 1.2.
+2. Proof of Theorem 1.2 for DVRs
 Consider the hypothesis below.
 Hypothesis 2.1.
-Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is a field).
-Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$ with $|G|=r$ invertible in $\mathcal{O}/\pi\mathcal{O}$.
+Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$).
+Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$.
+Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $\mathcal{O}$).
+Assume that $|G|$ is invertible in $\mathcal{O}$ i.e.
+$|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
+In this section, we will prove the following.
+Theorem 2.2.
+Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by psuedo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
 We can easily generalize Proposition 3.5 in [3] (which was for $n=2$) for any $n\in\mathbb{N}$ for our case. We provide the proof below for the sake of completeness.
-Proposition $2.$
+Proposition 2.3.
 Assume Hypothesis 2.1. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is injective.
@@
 So $YU=0$.
-The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is rIrI which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
+The image of $U$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is $rI$ which is invertible in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ since $r$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
 Consider the map $U:\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$.
@@
 Follows from Proposition 2.3 and Lemma 2.4. ∎
-Remark 2.3.1.
-Our proof of Lemma 2.3 fails when $\mathcal{O}$ has dimension $\geq 2$. Suppose
-$\mathcal{O}=k[X,Y]_{(X,Y)}$.
-Let $v=e_{1}/x^{2}+e_{2}/y^{2}\in K^{2}$ be such that $\sigma(v)=\lambda v$.
-Then note that for no choice of $t\in\mathcal{O}$ does $tv\in\mathcal{O}^{2}\setminus\mathfrak{m}\mathcal{O}^{2}$.
-So our induction fails at the base case $n=2$.
-3. Proof of Theorem 1.2
-Firstly we observe the following.
-Hypothesis 3.1.
-Consider the case in Hypothesis $1.$ Let $P$ be a maximal ideal of AA.
-Set $(\mathcal{O},(\pi))=(A_{P},PA_{P})$.
-Now $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}=K$.
-Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(\mathcal{O})$.
-Further since $|G|$ is invertible in AA, we have $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$.
-Thus for proving Theorem 1.2, it suffices to prove that $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
-Now let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$.
-Now we give the proof of Theorem 1.2 by proving $S$ is regular in Hypothesis 3.1.
+Now we give the proof of Theorem 2.2.
 Proof.
-Consider the short exact sequence
+Let $R=\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$. Consider the short exact sequence
 $0\rightarrow R\xrightarrow{\pi}R\rightarrow R/\pi R\rightarrow 0.$
@@
 $0\rightarrow R^{G}\xrightarrow{\pi}R^{G}\rightarrow(R/\pi R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{\pi}H^{1}(G,R)\rightarrow H^{1}(G,R/\pi R)\rightarrow\dots$
-Let $K=ker(\pi)=Img(\delta)$ in the above sequence.
-Now we have that $K$ is $R^{G}/\pi R^{G}$-module.
 By Theorem 6.5.8 in [4] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$.
-Hence $r\cdot K=0$.
-As $r$ is a unit in $R^{G}/\pi R^{G}$, we have $K=0$.
+As $r$ is a unit in $\mathcal{O}=(R^{G})_{0}$, we have that it is a unit in $R^{G}$.
+Thus $H^{1}(G,R)=0$.
 Hence we have the short exact sequence
@@
 Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections.
-Thus by Theorem 7.2.1 in [1], $(R/\pi R)^{G}$ is regular.
+Thus by Theorem 1.1, $(R/\pi R)^{G}$ is regular.
 Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$.
@@
 Thus $R^{G}/\pi R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
+3. Proof of Theorem 1.2 for regular local rings
+Consider the hypothesis below.
+Hypothesis 3.1.
+Let $(A,\mathfrak{m})$ be a regular local ring and let $K$ be its field of fractions.
+Consider a finite group $G\subseteq GL_{n}(A)$.
+Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$).
+Assume that $|G|$ is invertible in $A$.
+In this section, we will prove the following.
+Theorem 3.2.
+Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+Now consider the following.
+Hypothesis 3.3.
+Consider the case in Hypothesis 3.1.
+Let $d=\dim(A)$.
+Now Theorem 3.2 for the case $d=0$ (i.e.
+$A$ is a field) follows from Theorem 1.1.
+Thus we can assume $d\geq 1$.
+Let $\mathfrak{m}=(x_{1},\dots,x_{d})$.
+Set $x=x_{1}$ and let $\mathfrak{p}$ be the prime ideal $(x)$.
+We have an analogous result to Proposition 2.3 for regular local rings. Essentially the same proof goes through but we provide it for the sake of completeness.
+Lemma 3.4.
+Assume Hypothesis 3.3. For $n\in\mathbb{N}$, the natural map $\eta:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective.
+Proof.
+We consider elements of $GL_{n}(A)$ and $GL_{n}(A/\mathfrak{p})$ as matrices.
+Let $W\in G$ with $\eta(W)=I$.
+Then $W=I+Y$ where all entries of $Y$ are in $\mathfrak{p}$.
+Now $(I+Y)^{r}=W^{r}=I$ where $r=|G|$.
+Thus
+$rY+{r\choose 2}Y^{2}+\dots+\ {r\choose{r-1}}Y^{r-1}+Y^{r}=0.$
+Set
+$U=rI+\ {r\choose 2}Y+\dots+\ {r\choose{r-1}}Y^{r-2}+Y^{r-1}.$
+So $YU=0$.
+The image of $U$ in $GL_{n}(A/\mathfrak{p})$ is $rI$ which is invertible in $GL_{n}(A/\mathfrak{p})$ since $r$ is invertible in $A/\mathfrak{p}$.
+Consider the map $U:A^{n}\rightarrow A^{n}$.
+By Nakayama’s Lemma it follows that this map is surjective.
+Since $A^{n}$ is noetherian module over $A$, the map is an isomorphism.
+Thus $U\in GL_{n}(A)$.
+As $YU=0$, we get $Y=0$.
+So $\eta$ is injective.∎
+Now we give the proof of Theorem 3.2 with notations as in Hypothesis 3.3.
+Proof.
+We prove the result by induction on $d\geq 1$.
+Now the result for the case $d=1$ (i.e.
+$A$ is a DVR) follows from Theorem 2.2.
+So assume $d\geq 2$ and that the result is known for all regular local rings of dimension d1d-1.
+As $A$ is a regular local ring we have that $A_{\mathfrak{p}}$ is a regular local ring of dimension $1$, i.e.
+a DVR which has $K$ as its field of fractions.
+Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$.
+Now $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections.
+Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections.
+Note that we have a commutative diagram of ring homomorphisms,
+[…]
+where $\mu$ is the standard ring homomorphism of $A/\mathfrak{p}$ to its field of fractions which is injective.
+As we have the above commutative diagram, we also have the below commutative diagram of group homomorphisms with $\mu$ being injective.
+$GL_{n}(A)$$GL_{n}(A/\mathfrak{p})$$GL_{n}(A_{\mathfrak{p}})$$GL_{n}(\kappa(\mathfrak{p}))$$\scriptstyle{\iota}$$\scriptstyle{\gamma}$$\scriptstyle{\mu}$$\scriptstyle{\gamma^{\prime}}$
+We have that $\gamma|_{G}:G\rightarrow GL_{n}(A/\mathfrak{p})$ is injective by Lemma 3.4.
+Thus we have that $\mu\circ\gamma|_{G}$ is injective.
+Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$.
+So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections.
+By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ to be regular.
+Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$.
+Consider the short exact sequence
+$0\rightarrow R\xrightarrow{x}R\rightarrow R/\mathfrak{p}R\rightarrow 0.$
+We have the long exact sequence
+$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\xrightarrow{\delta}H^{1}(G,R)\xrightarrow{x}H^{1}(G,R)\rightarrow H^{1}(G,R/\mathfrak{p}R)\rightarrow\dots$
+By Theorem 6.5.8 in [4] we have $r\cdot H^{1}(G,R)=0$ where $r=|G|$.
+As $r$ is a unit in $A=(R^{G})_{0}$, we have that it is a unit in $R^{G}$.
+Thus $H^{1}(G,R)=0$.
+Hence we have the short exact sequence
+$0\rightarrow R^{G}\xrightarrow{x}R^{G}\rightarrow(R/\mathfrak{p}R)^{G}\rightarrow 0.$
+Thus $R^{G}/\mathfrak{p}R^{G}\cong(R/\mathfrak{p}R)^{G}$ which is regular.
+Let $\mathfrak{M}$ be the maximal homogenous ideal of $R^{G}$.
+Since $x\not\in\mathfrak{M}^{2}$.
+Thus $R^{G}/\mathfrak{p}R^{G}$ being regular implies that $S=R^{G}$ is regular.∎
+4. Proof of Theorem 1.2 for regular domains
+Now we complete the proof of Theorem 1.2.
+Proof.
+Now $A$ is a regular domain and $K$ is its field of fractions.
+Let $P$ be a maximal ideal of $A$.
+So $(A_{P},PA_{P})$ is a regular local ring with $K$ as the field of fractions.
+Also, as $G$ is a subgroup of $GL_{n}(A)$, it is naturally a subgroup of $GL_{n}(A_{P})$.
+Further since $|G|$ is invertible in $A$, we have $|G|$ is invertible in $A_{P}$.
+Thus for proving Theorem 1.2, it suffices to prove that $(A_{P}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular which follows from Theorem 3.2.
+∎
+Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius.

### v3

@@
 Assume that $|G|$ is invertible in $A$.
-We prove that if $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+We prove that $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
 Key words and phrases:
@@
 1. Introduction
+We begin by recalling the notion of pseudo-reflections.
+Definition 1.1 (§7.1, [1]).
+Let $K$ be a field. An element $\sigma\in GL_{n}(K)$ is called a pseudo-reflection if $ord(\sigma)$ is finite and $dim(ker(\sigma-1))=n-1$, i.e. whose fixed points have codimension one.
 We have the famous Chevalley-Shephard-Todd Theorem (Theorem 7.2.1 in [1]) which states:
@@
 Assume that $|G|$ is invertible in $K$.
-We have $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
+We have $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over $K$.
 Mundelius had generalized Theorem 1.2 to the rings of invariants of pseudo-reflection groups over Dedekind domains (Theorem 3.8 in [3]).
@@
 Assume that $|G|$ is invertible in $A$.
-If $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections, then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+We have $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections if and only if $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+One direction of the above theorem follows from Theorem 1.2. Here’s the proof.
+Proof.
+Suppose $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+Now $G$ acts linearly on the ring $A[X_{1},X_{2},\dots,X_{n}]$ (fixing $A$) and $|G|$ is invertible in $A$.
+As $K$ is the field of fractions of $A$, so $G$ also acts linearly on the ring $K[X_{1},X_{2},\dots,X_{n}]$ (fixing $K$) and $|G|$ is invertible in $K$.
+Now as $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a localization of $(A[X_{1},X_{2},\dots,X_{n}])^{G}$, we have that $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+Therefore by Theorem 5.3.3 in [4] and by Theorem 19.2 in [2], $(K[X_{1},X_{2},\dots,X_{n}])^{G}$ is a polynomial ring over the field $K$.
+Hence by Theorem 1.2, $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections .
+∎
+The subsequent sections of the paper are dedicated to proving the other direction.
 In Section 2, we prove Theorem 1.3 for the case when $A$ is a DVR as Theorem 2.2.
@@
 Hypothesis 2.1.
-Let $(\mathcal{O},(\pi))$ be a DVR and $\mathcal{O}_{\pi}$ be the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$).
+Assume that $(\mathcal{O},(\pi))$ is a DVR and $\mathcal{O}_{\pi}$ is the localization of $\mathcal{O}$ at the multiplicative set $\{\pi^{i}\}_{i}$ (which is the field of fractions of $\mathcal{O}$).
 Consider a finite group $G\subseteq GL_{n}(\mathcal{O})$.
-Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $\mathcal{O}$).
+Consider the ring $\mathcal{O}[X_{1},X_{2},\dots,X_{n}]$ and assume that $G$ acts linearly on the ring (fixing $\mathcal{O}$).
 Assume that $|G|$ is invertible in $\mathcal{O}$ i.e.
@@
 Theorem 2.2.
-Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by psuedo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+Assume Hypothesis 2.1. Consider $G\subseteq GL_{n}(\mathcal{O}_{\pi})$ is generated by pseudo-reflections. Then $(\mathcal{O}[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
 We can easily generalize Proposition 3.5 in [5] (which was for $n=2$) for any $n\in\mathbb{N}$ for our case. We provide the proof below for the sake of completeness.
@@
 Lemma 2.4.
-Assume Hypothesis 2.1.
-Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$.
-Then for $n\geq 2$ and any psuedo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
+Assume Hypothesis 2.1. Suppose that $G$ is generated by pseudo-reflections in
+$GL_{n}(\mathcal{O}_{\pi})$.
+Then for $n\geq 2$ and any pseudo-reflection $\sigma\in G\subseteq GL_{n}(\mathcal{O}_{\pi})$, there exists a basis $\{w_{1},w_{2},\dots w_{n}\}$ of $\mathcal{O}^{n}$ with $\sigma(w_{i})=w_{i}$ for $1\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}$ where $\lambda$ is an $m$-th root of unity where $m=ord(\sigma)$.
 Proof.
-We prove the result by induction on $n\geq 2$.
-Let $n=2$ and let $\{v_{1},v_{2}\}$ be a basis of $\mathcal{O_{\pi}}^{2}$ such that $\sigma(v_{1})=v_{1}$ and $\sigma(v_{2})=\lambda v_{2}$.
-We can assume $v_{1},v_{2}\in\mathcal{O}^{2}\backslash\pi\mathcal{O}^{2}$.
-Note $\lambda\neq 1$.
-Let $w_{1}=v_{1}$ and let $\{w_{1},w_{2}\}$ be a basis of $\mathcal{O}^{2}$.
-Let $v_{2}=\alpha w_{1}+\beta w_{2}$.
-Note that $\beta\neq 0$.
-Now $\sigma(v_{2})=\lambda v_{2}=\lambda\alpha w_{1}+\lambda\beta w_{2}$.
-Also we have $\sigma(v_{2})=\alpha\sigma(w_{1})+\beta\sigma(w_{2})=\alpha w_{1}+\beta\sigma(w_{2})$.
-Thus comparing both equations we have
-(2.1) $\beta\sigma(w_{2})=(\lambda-1)\alpha w_{1}+\lambda\beta w_{2}$
+Let $\{v_{1},\ldots v_{n}\}$ be a basis of $(\mathcal{O}_{\pi})^{n}$ such that $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$.
 Let $\lambda>$ be the subgroup of $\mathcal{O}^{\times}$ generated by $\lambda$.
-Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid r$.
+Since $ord(\lambda)=ord(\sigma)$, we have $|ord(\lambda)|\mid|G|$.
 As $|G|$ is invertible in $\mathcal{O}/\pi\mathcal{O}$, so $|<\lambda>|$ is also invertible in $\mathcal{O}/\pi\mathcal{O}$.
@@
 So $\lambda-1\in\mathcal{O}^{\times}$.
-We claim that $\beta\in\mathcal{O}^{\times}$.
-Assume on the contrary $\beta\in\pi\mathcal{O}$.
-Since $\alpha w_{1}+\beta w_{2}=v_{2}\not\in\pi\mathcal{O}^{2}$.
-Thus $\alpha\not\in\pi\mathcal{O}$.
-Going modulo $\pi$ in the Equation 2.1, we have $(\bar{\lambda}-1)\bar{\alpha}\bar{w_{1}}=0$ which is a contradiction.
-Thus we can write Equation 2.1 as $\sigma(w_{2})=\lambda w_{2}+\theta w_{1}$ for some $\theta\in\mathcal{O}$.
-Then letting $w^{\prime}_{2}=\frac{\theta}{\lambda-1}w_{1}+w_{2}$, we have that $\{w_{1},w^{\prime}_{2}\}$ is a basis and $\sigma(w^{\prime}_{2})=\lambda w^{\prime}_{2}$.
-Let $n\geq 3$ and $\{v_{1},v_{2},\dots,v_{n}\}$ be a basis of $\mathcal{O}_{\pi}$ with $v_{i}\in\mathcal{O}^{n}\backslash\pi\mathcal{O}^{n}$ and $\sigma(v_{i})=v_{i}$ for $1\leq i\leq n-1$ and $\sigma(v_{n})=\lambda v_{n}$ and $\lambda\neq 1$.
-Let $E=\mathcal{O}v_{1}$ and $w_{1}=v_{1}$.
-We have the following exact sequence.
-$0\rightarrow E\rightarrow W=\mathcal{O}^{n}\rightarrow\bar{W}=\mathcal{O}^{n-1}\rightarrow 0.$
-Now $\sigma|_{\bar{W}}$ is also a psuedo-reflection.
-By induction we have a basis $\{\bar{w_{2}},\bar{w_{3}},\dots,\bar{w_{n}}\}$ of $\bar{W}$ with $\sigma(\bar{w_{i}})=\bar{w_{i}}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda\bar{w_{n}}$.
-Now $\{w_{1},w_{2},\dots,w_{n}\}$ is a basis of $W$ and $\sigma(w_{1})=w_{1}$ and $\sigma(w_{i})=w_{i}+a_{i}w_{1}$ for $2\leq i\leq n-1$ and $\sigma(w_{n})=\lambda w_{n}+a_{n}w_{1}$.
-We will show that $a_{2}=a_{3}=\dots=a_{n-1}=0$.
-Consider $2\leq i\leq n-1$.
-Now $\sigma(w_{i})=w_{i}+a_{i}w_{1}$.
-Thus $\sigma^{j}(w_{i})=w_{i}+ja_{i}w_{1}$.
-Now $\sigma^{m}=id$ where $m=ord(\sigma)=ord(\lambda)$.
-Thus $w_{i}=w_{i}+ma_{i}w_{1}$.
-Hence $ma_{i}w_{1}=0$.
-As $m$ is invertible in $\mathcal{O}$, we have $a_{i}=0$.
-Also by Proposition 2.3, $\lambda-1$ is invertible in $\mathcal{O}$.
-Thus $w_{1},w_{2},\dots,w_{n-1},w^{\prime}_{n}$ is the required basis where $w^{\prime}_{n}=\frac{a_{n}}{\lambda-1}w_{1}+w_{n}$ and $\sigma(w^{\prime}_{n})=\lambda w^{\prime}_{n}$.
+Consider the endomorphism $\epsilon=(\lambda-1)^{-1}(\sigma-id)\colon\mathcal{O}^{n}\rightarrow\mathcal{O}^{n}$.
+By considering $\epsilon_{\pi}$ it follows that $\epsilon$ is an idempotent.
+So $\mathcal{O}^{n}=\ker\epsilon\oplus\ker(1-\epsilon)$.
+We note that $\operatorname{rank}(\ker(\epsilon))=n-1$ and $\operatorname{rank}(\ker(1-\epsilon))=1$.
+On $\ker(\epsilon)$ the map $\sigma$ acts as identity while on $\ker(1-\epsilon)$ the map $\sigma$ acts as multiplication by $\lambda$.
+The result follows.
 ∎
@@
 Assume Hypothesis 2.1.
-Suppose that $G$ is generated by psuedo-reflections in $\mathcal{O}_{\pi}$.
-Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by psuedo-reflections.
+Suppose that $G$ is generated by pseudo-reflections in $GL_{n}(\mathcal{O}_{\pi})$.
+Then for $n\geq 2$, the image of $G$ in $GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is also generated by pseudo-reflections.
 Proof.
@@
 Now $R/\pi R=(\mathcal{O}/\pi\mathcal{O})[X_{1},X_{2},\dots,X_{n}]$ and $\mathcal{O}/\pi\mathcal{O}$ is a field.
-Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by psuedo-reflections.
+Also by Corollary 2.4.1 we have that $G\subseteq GL_{n}(\mathcal{O}/\pi\mathcal{O})$ is generated by pseudo-reflections.
 Thus by Theorem 1.2, $(R/\pi R)^{G}$ is regular.
@@
 Hypothesis 3.1.
-Let $(A,\mathfrak{m})$ be a regular local ring and let $K$ be its field of fractions.
+Assume that $(A,\mathfrak{m})$ is a regular local ring and $K$ is its field of fractions.
 Consider a finite group $G\subseteq GL_{n}(A)$.
-Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and let $G$ act linearly on the ring (fixing $A$).
+Consider the ring $A[X_{1},X_{2},\dots,X_{n}]$ and assume that $G$ acts linearly on the ring (fixing $A$).
 Assume that $|G|$ is invertible in $A$.
@@
 Theorem 3.2.
-Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
+Assume Hypothesis 3.1. Consider $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections. Then $(A[X_{1},X_{2},\dots,X_{n}])^{G}$ is regular.
 Now consider the following.
 Hypothesis 3.3.
-Consider the case in Hypothesis 3.1.
+Consider the assumptions in Hypothesis 3.1.
 Let $d=\dim(A)$.
@@
 Let $\kappa(\mathfrak{p})=A_{\mathfrak{p}}/\mathfrak{p}A_{\mathfrak{p}}$ which is the field of fractions of $A/\mathfrak{p}$.
-Now $G\subseteq GL_{n}(K)$ is generated by psuedo-reflections.
-Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections.
+Now $G\subseteq GL_{n}(K)$ is generated by pseudo-reflections.
+Thus by Corollary 2.4.1, we have that the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by pseudo-reflections.
 Note that we have a commutative diagram of ring homomorphisms,
@@
 Now $\gamma^{\prime}\circ\iota|_{G}=\mu\circ\gamma|_{G}$.
-So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by psuedo-reflections.
-By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ to be regular.
+So the image of $G$ in $GL_{n}(\kappa(\mathfrak{p}))$ is generated by pseudo-reflections.
+By induction we have $(A/\mathfrak{p}\ [X_{1},X_{2},\dots X_{n}])^{G}$ is regular.
 Let $R=A[X_{1},X_{2},\dots,X_{n}]$ and $S=R^{G}$.
@@
 ∎
-Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius.
+5. Conflict of interest statement
+The authors have no conflict of interest to declare that are relevant to this article.
+6. Data availability statement
+The manuscript has no associated data.
+Acknowledgements: We would like to thank Prof. Gregor Kemper for making us aware of the work of David Mundelius. We also thank the two referees for many pertinent comments.